### v1

# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

Abstract
Counterfactual image generation aims to simulate realistic visual outcomes under specific causal interventions. Diffusion models have recently emerged as a powerful tool for this task, combining DDIM inversion with conditional generation via classifier-free guidance (CFG). However, standard CFG applies a single global weight across all conditioning variables, which can lead to poor identity preservation and spurious attribute changes—a phenomenon known as attribute amplification. To address this, we propose Decoupled Classifier-Free Guidance (DCFG), a flexible and model-agnostic framework that introduces group-wise conditioning control. DCFG builds on an attribute-split embedding strategy that disentangles semantic inputs, enabling selective guidance on user-defined attribute groups. For counterfactual generation, we partition attributes into intervened and invariant sets based on a causal graph and apply distinct guidance to each. Experiments on CelebA-HQ, MIMIC-CXR, and EMBED show that DCFG improves intervention fidelity, mitigates unintended changes, and enhances reversibility, enabling more faithful and interpretable counterfactual image generation.
1 Introduction
Counterfactual generation is fundamental to causal reasoning [37, 40, 1], allowing us to explore hypothetical scenarios, such as How would this patient’s disease have progressed if treatment A had been administered instead of treatment B? The ability to answer such causal questions is important across various domains, such as healthcare [46], fairness [26, 65] and scientific discovery [33]. There has been a growing interest in generating counterfactual images using deep generative models [23, 67], aiming to simulate how visual data would change under hypothetical interventions. Recent works extend the framework of Structural Causal Models (SCMs) [38, 40, 2] with deep generative models, performing counterfactual generation via a three-step procedure: abduction-action-prediction [36, 9, 62, 61, 8, 47]. These works typically parameterise their image-generating mechanisms with normalising flows [35, 59], GANs [14], VAEs [22], and HVAEs [5, 56]. However, these generative backbones often exhibit trade-offs between fidelity and flexibility [43, 64]. In this work, we explore diffusion models as a more powerful alternative for counterfactual inference.
Diffusion models [50, 54, 17] have emerged as the state-of-the-art approach for image synthesis, achieving unprecedented fidelity and perceptual quality [10, 41]. Recent works have utilised diffusion models for counterfactual generation [34, 24, 46, 45, 3, 4, 29, 58, 44, 12, 53, 39, 25]. A common approach involves DDIM inversion to encode observed images into latent representations, followed by conditional generation under modified attributes. Conditioning is typically enforced via discriminative score functions, either with external classifiers [10] or through classifier-free guidance [16]. This combination of DDIM inversion and conditional decoding has become a dominant paradigm in diffusion-based image editing [7, 31, 57, 15, 11]. Building upon this paradigm, we investigate its use for high-fidelity counterfactual inference via a novel classifier-free guidance strategy.
Classifier-Free Guidance (CFG) [16] has become the standard approach for conditioning diffusion models without requiring external classifiers. In counterfactual generation, CFG plays a crucial role in ensuring that the intended intervention is faithfully reflected in the output. While recent works have proposed refinements to CFG to enhance sample fidelity and controllability [6, 21, 27], standard CFG applies a single global guidance weight across all conditioning attributes, regardless of which ones are intended to change. This can lead to over-editing or unintended modifications of attributes that should remain fixed—a phenomenon known as attribute amplification [63], where invariant factors are spuriously exaggerated. While Xia et al. [63] observed this issue in HVAE-based counterfactual models [9] as a result of training-time entanglement, we show that a similar failure mode can arise at inference time in diffusion models due to the indiscriminate application of global guidance. Such behaviour not only violates the underlying causal assumptions—by modifying attributes outside the intervention—but can also cause the generation trajectory to drift away from the original data manifold, degrading identity preservation [66, 31, 55].
To address the limitations of standard CFG, we propose Decoupled Classifier-Free Guidance (DCFG), a model-agnostic method that requires no changes to the diffusion model or loss function, and only modifies the conditioning embedding. DCFG can be readily applied to any diffusion model employing CFG. A key ingredient in our approach is an attribute-split conditioning embedding strategy, which disentangles semantic attributes in the embedding space and enables selective masking and group-wise modulation at inference time. Unlike standard CFG—which applies a single global guidance weight across all conditioning attributes—DCFG assigns separate weights to attribute groups, allowing for fine-grained, interpretable control over the generative process. While conceptually related to compositional diffusion approaches, our method differs significantly: Shen et al. [49] apply pixel-wise spatial masks to modulate guidance locally, and Liu et al. [28] rely on multiple conditional diffusion models fused via shared score functions. In contrast, DCFG uses a single model and modulates guidance at the semantic attribute level. For counterfactual generation, we instantiate DCFG by partitioning attributes into intervened and invariant sets based on their roles in a structural causal graph, and apply distinct guidance to each group. Crucially, by decoupling guidance and focusing it solely on the intended intervention, DCFG reduces the risk of the generation trajectory drifting away from the original data manifold [66, 31, 55]. While our experiments focus on causal groupings, the DCFG framework is general and supports arbitrary partitions of semantic attributes. The contributions of this paper are:
  • We introduce an attribute-split embedding strategy that disentangles semantic conditioning variables in the embedding space, enabling selective masking and modular guidance for DCFG.
  • We propose Decoupled Classifier-Free Guidance (DCFG), a simple and flexible extension of CFG that assigns separate guidance weights to groups of attributes. DCFG is model-agnostic, training-compatible with existing pipelines, and supports arbitrary attribute groupings at inference time.
  • We instantiate DCFG for counterfactual image generation by grouping attributes into intervened and invariant sets based on a structural causal graph, and show improved alignment with intervention targets while preserving irrelevant features.
2 Background
2.1 Causality
Structural Causal Models. SCMs [37] consist of a triplet $\langle U,A,F\rangle$, where $U=\{u_{i}\}_{i=1}^{K}$ denotes the set of exogenous (latent) variables, $A=\{a_{i}\}_{i=1}^{K}$ the set of endogenous (observed) variables, and $F=\{f_{i}\}_{i=1}^{K}$ a collection of structural assignments such that each variable $a_{k}$ is determined by a function $f_{k}$ of its parents $\mathbf{pa}_{k}\subseteq A\setminus a_{k}$ and its corresponding noise $u_{k}$: $a_{k}\coloneqq f_{k}(\mathbf{pa}_{k},u_{k})$. SCMs enable causal reasoning and interventions via the do-operator, e.g., setting a variable $a_{k}$ to a fixed value $c$ through $do(a_{k}\coloneqq c)$. In this work, we focus on generating image counterfactuals and implement the underlying image synthesis mechanism using diffusion models.
Counterfactual inference. We define an image as $\mathbf{x}$, which is generated from its causal parents $\mathbf{pa}$ and an exogenous noise variable $\mathbf{u}$ via a structural assignment: $\mathbf{x}:=f(\mathbf{u},\mathbf{pa})$. Counterfactual reasoning proceeds in three conceptual steps [37] - (1) Abduction: infer the latent noise $\mathbf{u}$ from the observed data and its parents: $\mathbf{u}:=f^{-1}(\mathbf{x},\mathbf{pa})$; (2) Action: apply an intervention to alter selected parent variables, yielding $\widetilde{\mathbf{pa}}$; (3) Prediction: propagate the effect of the intervention through the model to compute the counterfactual $\tilde{\mathbf{x}}:=f(\mathbf{u},\widetilde{\mathbf{pa}})$. Recent advancements have sought to implement these steps using deep learning models, with VAE [36], HVAE [9], diffusion models [44, 45, 24, 12]. The general idea is to model the structural assignment $f_{\theta}$ and its inverse $f^{-1}_{\theta}$ using neural networks with trainable parameters $\theta$.
2.2 Diffusion Models
DMs [51, 17] are latent variable models which are trained to generate images by gradually removing Gaussian noise from $x_{T}\sim N(0,I)$ over $T$ timesteps. Given a clean sample $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$, the forward process at timestep $t$ follows:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\,\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\{\alpha_{t}\}_{t=1}^{T}$ is a variance schedule defined over $t\in[1,T]$.
To learn the reverse process, a parameterized network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ is trained to predict the added noise from noisy inputs. We adopt the conditional diffusion model formulation, where $\mathbf{c}$ denotes the embedding of semantic parent attributes $\mathbf{pa}$ used as conditioning input. The training objective minimizes the noise prediction loss:
[…]
At inference time, data samples are generated by progressively denoising $\mathbf{x}_{t}$. Following Eq (12) in Song et al. [52], the denoising step from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ is:
$\mathbf{x}_{t-1}=\underbrace{\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})}{\sqrt{\alpha_{t}}}\right)}_{\text{predicted }\mathbf{x}_{0}\text{}}+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})}_{\text{direction to }\mathbf{x}_{t}\text{}}+\underbrace{\sigma_{t}\boldsymbol{\epsilon}_{t}}_{\text{noise}},$ (3)
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Setting $\sigma_{t}=0$ yields a deterministic sampling process known as DDIM [52], which defines an invertible trajectory between data and latent space. Following [24, 44, 12, 39], we adopt this DDIM formulation for counterfactual generation, as detailed below.
Abduction. We implement the abduction function $\mathbf{u}:=f^{-1}_{\theta}(\mathbf{x},\mathbf{pa})$ using the DDIM forward trajectory. Given an observed image $\mathbf{x}_{0}$ (i.e. $\mathbf{x}$ in $f^{-1}_{\theta}$) and conditioning vector $\mathbf{c}$ (the embedding of semantic parents $\mathbf{pa}$), the latent $\mathbf{x}_{T}$ serves as a deterministic estimate of the exogenous noise $\mathbf{u}$:
$\mathbf{x}_{t+1}=\sqrt{{\alpha}_{t+1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t+1}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c}),\quad t=0,\dots,T-1,$ (4)
where the clean estimate $\hat{\mathbf{x}}_{0}$ at each step is computed as:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-{\alpha}_{t}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right),$ (5)
Action. We apply an intervention to the semantic attributes $\mathbf{pa}$ (e.g., do(Male = 1)), and propagate the effect through the causal graph using invertible flows as in [36, 9]. This yields the counterfactual attribute vector $\widetilde{\mathbf{pa}}$ and its embedding $\tilde{\mathbf{c}}$.
Prediction. We implement the structural assignment $\tilde{\mathbf{x}}:=f_{\theta}(\mathbf{u},\widetilde{\mathbf{pa}})$ under the modified condition $\tilde{\mathbf{c}}$ using the DDIM reverse trajectory, with $\mathbf{u}=\mathbf{x}_{T}$ the exogenous noise estimated in eq. 4:
$\mathbf{x}_{t-1}=\sqrt{{\alpha}_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t-1}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}),\quad t=T,\dots,1,$ (6)
where $\hat{\mathbf{x}}_{0}$ is computed as in eq. 5. The final output $\tilde{\mathbf{x}}_{0}$ is the predicted counterfactual $\tilde{\mathbf{x}}$.
In practice, decoding under the counterfactual condition $\tilde{\mathbf{c}}$ using the conditional denoiser alone may be insufficient for producing effective counterfactuals. Additional guidance is often required to steer generation toward the desired intervention [44, 45, 24, 12, 58, 53, 39, 25]. In this work, we adopt classifier-free guidance (CFG) to enhance counterfactual fidelity and alignment with the specified intervention.
2.3 Classifier-Free Guidance
CFG [16] is a widely adopted technique in conditional diffusion models. It enables conditional generation without requiring an external classifier by training a single denoising model to operate in both conditional and unconditional modes. During training, the model learns both $p_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})$ and $p_{\theta}(\mathbf{x}_{t}\mid\varnothing)$ by randomly replacing $\mathbf{c}$ with a null token $\varnothing$.
At inference time, CFG samples from a sharpened posterior:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{c})\propto p(\mathbf{x}_{t})\cdot p(\mathbf{c}\mid\mathbf{x}_{t})^{\omega},$ (7)
where $\omega\geq 0$ is a guidance weight controlling the strength of conditioning. This corresponds to interpolating between the unconditional and conditional scores:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{c})=\nabla\log p(\mathbf{x}_{t})+\omega\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{c})-\nabla\log p(\mathbf{x}_{t})\right).$ (8)
In practice, this is implemented by interpolating denoised predictions:
$\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\mathbf{c})=\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t\mid\varnothing)+\omega\cdot\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)\right).$ (9)
With CFG, abduction is the same as in eq. 4, and action remains unchanged. The only difference lies in prediction, where the conditional denoiser is replaced with the guided score $\boldsymbol{\epsilon}_{\mathrm{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})$ to enhance counterfactual effectiveness [45, 24]:
$\hat{\mathbf{x}}_{0}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})\right),$ (10)
$\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t-1}}\cdot\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}).$ (11)
Despite its effectiveness, standard CFG applies a single global guidance weight $\omega$ uniformly across the entire counterfactual embedding $\tilde{\mathbf{c}}$, which typically encodes multiple attributes—some of which may not have been altered during the intervention. In counterfactual generation, however, only a subset of attributes in $\tilde{\mathbf{c}}$ (i.e., those affected by the intervention) should be emphasized, while the remaining attributes should remain invariant. Applying the same guidance strength to all elements of $\tilde{\mathbf{c}}$ violates this principle, and can cause unintended changes to invariant attributes. This misalignment is called attribute amplification [63], which violates the causal relationship as pre-defined in structural causal models and undermines counterfactual soundness [32].
To address these limitations, we propose a structured alternative that assigns separate guidance weights to semantically or causally defined groups of attributes.
3 Method
In this section, we present our Decoupled Classifier-Free Guidance (DCFG) for counterfactual image generation. We begin by introducing an attribute-split conditioning embedder (section 3.1), which explicitly separates individual attributes in the embedding space to enable selective control. Building on this, we then describe our DCFG formulation (section 3.2), which allows distinct guidance strengths to be applied to different subsets of attributes. Finally, we present how DCFG is integrated into DDIM-based counterfactual inference (section 3.3), detailing its application across abduction, action, and prediction steps using causally defined attribute groupings.
3.1 Attribute-Split Conditioning Embedding
In practice, raw conditioning inputs such as discrete labels or structured attributes (e.g., a patient’s sex, race, disease status) are not used directly in diffusion models but transformed into dense vectors using embedding functions—typically via small MLPs [10], convolutional encoders [68], or transformer-based text encoders [16, 42]. These embeddings align semantic or categorical inputs with the model’s internal representation space, but conventional designs often entangle multiple attributes into a single vector, making it difficult to independently control them during sampling.
To address this, we introduce an attribute-split conditioning embedding that preserves the identity of each attribute in the embedding space. Let $pa_{i}$ denote the raw value of the $i$-th parent attribute (e.g., a binary indicator or scalar). Each $pa_{i}$ is embedded independently via a dedicated MLP $\mathcal{E}_{i}:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d}$, and the final condition vector is formed by concatenating the outputs:
$\mathbf{c}=\operatorname{concat}\left(\mathcal{E}_{1}(pa_{1}),\dots,\mathcal{E}_{K}(pa_{K})\right),\quad\mathbf{c}\in\mathbb{R}^{K\cdot d}.$ (12)
This architecture provides a flexible representation where each attribute is explicitly disentangled at the embedding level. As a result, we can selectively null-tokenize or modulate individual attributes at inference time—enabling fine-grained control. Throughout the rest of the paper, we denote the semantic attribute vector as $\mathbf{pa}$ and the corresponding embedding vector as $\mathbf{c}$, as defined in eq. 12.
3.2 Decoupled Classifier-Free Guidance
To overcome the limitations of CFG and enable more precise, causally aligned control in counterfactual image generation, we propose Decoupled Classifier-Free Guidance (DCFG). Rather than applying a single guidance weight uniformly to the entire conditioning vector, we partition semantic attributes $\mathbf{pa}$ into $M$ disjoint groups $\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}$, and apply a separate guidance weight $\omega_{m}$ to each group. Let $\mathbf{pa}=(pa_{1},\dots,pa_{K})$ denote the vector of semantic parent attributes. Under the assumption that the groups $\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}$ are conditionally independent given the latent variable $\mathbf{x}_{t}$, we obtain the following factorized proxy posterior:
Proposition 1 (Proxy Posterior for DCFG).
Under the assumption $p(\mathbf{pa}\mid\mathbf{x}_{t})=\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})$, the sharpened proxy posterior becomes:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})\propto p(\mathbf{x}_{t})\cdot\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})^{\omega_{m}},$ (13)
where $\omega_{m}\geq 0$ controls the guidance strength for group $m$.
A full derivation and gradient-based justification for this proxy posterior are provided in Appendix B. The corresponding gradient used in score-based diffusion sampling is:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})=\nabla\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla\log p(\mathbf{x}_{t})\right).$ (14)
[…]
The final guided score used in the diffusion model is computed as:
[…]
The proposed DCFG framework is highly flexible, as it allows arbitrary groupings of attributes, regardless of whether attributes within a group are mutually independent or not. The only assumption required is that different groups are conditionally independent given the latent variable $\mathbf{x}_{t}$. This flexibility enables a wide range of configurations. For instance, setting $M=1$ recovers standard global CFG, while increasing $M$ provides progressively finer-grained control, including per-attribute guidance ($M=K$) as an extreme case where we assume all attributes are independent of each other.
3.3 DCFG for Counterfactual Generation
We now detail how DCFG is integrated into DDIM-based counterfactual inference.
Abduction. The abduction step proceeds as in eq. 4, where the conditioning vector $\mathbf{c}$ is obtained by embedding the semantic parent attributes $\mathbf{pa}$ using the attribute-split encoder defined in eq. 12.
Action. As in previous setups, we apply a causal intervention to obtain a modified semantic vector $\widetilde{\mathbf{pa}}$. This is then embedded into the counterfactual conditioning vector $\tilde{\mathbf{c}}$ via the attribute-split embedder:
$\tilde{\mathbf{c}}=\operatorname{concat}\left(\mathcal{E}_{1}(\widetilde{pa}_{1}),\dots,\mathcal{E}_{K}(\widetilde{pa}_{K})\right).$ (17)
Prediction. The prediction step uses the DCFG-guided reverse DDIM trajectory:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-{\alpha}_{t}}\cdot\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})\right),$ (18)
$\mathbf{x}_{t-1}=\sqrt{{\alpha}_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t-1}}\cdot\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}),$ (19)
where $\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})$ is computed as in eq. 16 using counterfactual conditioning embedding $\tilde{\mathbf{c}}$.
For the counterfactual generation task, we follow a two-group partitioning of attributes based on the causal graph. The affected group $\mathbf{pa}^{\text{aff}}$ consists of attributes that are directly intervened on, along with their causal descendants. The invariant group $\mathbf{pa}^{\text{inv}}$ comprises attributes that are not affected by the intervention and are expected to remain unchanged. These groups are assumed to be conditionally independent given the latent $\mathbf{x}_{t}$, consistent with the d-separation implied by the post-intervention causal graph. Under this setup, eq. 16 is instantiated with $M=2$ groups, applying separate guidance weights $\omega_{\text{aff}}$ and $\omega_{\text{inv}}$ to the affected and invariant groups, respectively. In practice, we find that $\omega_{\text{inv}}=1.0$ is typically sufficient when $\omega_{\text{aff}}$ is small. However, as $\omega_{\text{aff}}$ increases, the resulting intervention may introduce unintended drift in $\mathbf{pa}^{\text{inv}}$ due to more aggressive changes to the image. In such cases, raising $\omega_{\text{inv}}$ can help maintain the desired invariance. Determining the optimal $(\omega_{\text{aff}},\omega_{\text{inv}})$ pair remains an open problem for future exploration. Fig. A.8 visualizes the effect of $\omega_{\text{inv}}$.
4 Experiments
In this section, we demonstrate the benefits of the proposed approach across three public datasets. For each dataset, we train a diffusion model with the same architecture and training protocol, detailed in appendix C. We compare our DCFG against the standard CFG baseline. In all results, settings labeled as $\omega=X$ correspond to standard classifier-free guidance (CFG) with a global guidance weight. In contrast, configurations denoted by $\omega_{\text{aff}}=X$, $\omega_{\text{inv}}=Y$ represent our proposed DCFG, where separate guidance weights are applied to the intervened and invariant attribute groups, respectively. Following Melistas et al. [30], Monteiro et al. [32], we evaluate counterfactual quality using two metrics. Effectiveness ($\Delta$): Measured by a pretrained classifier as the change in AUROC for intervened attributes relative to $\omega=1.0$ (no CFG). Higher $\Delta$ indicates stronger intervention effect; large $\Delta$ on invariant attributes indicates unintended amplification. Reversibility: Assesses how well counterfactuals can be reversed to the original image using inverse interventions. We report MAE and LPIPS; lower values indicate better identity preservation. See section A.2 for details.
4.1 CelebA
Figure 1: Comparison of $\Delta$ metrics under different interventions in CelebA-HQ. Left: Intervention on Smiling. Right: Intervention on Young. Both use baseline $\omega=1.0$. Under global CFG, increasing $\omega$ boosts the intended attribute but amplifies non-target attributes. DCFG achieves similar improvements on the target attribute while mitigating amplification. See section D.2 for full quantitative results.
We begin our evaluation on the CelebA-HQ dataset [20], using Smiling, Male, and Young as independent binary parent attributes. Although some variables (e.g., Smiling and Mouth Open) may be causally linked in the real world, we deliberately adopt this simplified setup to isolate and diagnose inference-time failures caused by global CFG. Under this assumption, any unintended change in non-intervened attributes can be confidently attributed to attribute amplification by standard CFG, rather than valid downstream causal effects. Refer to section D.1 for more details.
Fig. 1 presents the $\Delta$ metrics under different guidance strategies for two separate interventions: do(Smiling) and do(Young). As the global guidance weight $\omega$ increases (left to right side of each plot), the $\Delta$ of the intervened attribute improves, but so do the $\Delta$ values of attributes that should remain invariant, indicating undesirable amplification. In contrast, the right side of each plot shows results for DCFG, where distinct weights are applied to affected ($\omega_{\text{aff}}$) and invariant ($\omega_{\text{inv}}$) attribute groups. This decoupled formulation achieves comparable or stronger improvement on the intervened attribute while keeping the others stable, validating DCFG ’s ability to produce more disentangled and effective counterfactuals.
Figure 2: Counterfactual generations in CelebA-HQ ($64\times 64$). Each row compares global CFG (left) and DCFG (right) across guidance weights. Top: global CFG causes amplification of Smiling under do(Male); Middle: do(Young) suppresses Male (i.e. amplifies $\texttt{Male}=no$); Bottom: do(Smiling) makes the subject appear older, adds glasses, and alters identity. DCFG mitigates these unintended changes and preserves non-intervened attributes. Refer to section D.3 for more visual results.
Fig. 2 illustrates how global CFG can introduce unitended changes by uniformly amplifying all conditioning signals, even when only one attribute is meant to change. In the top row, applying do(Male=no) with increasing $\omega$ inadvertently amplifies Smiling; in the middle row, do(Young=no) reduces Male expression; and in the bottom row, do(Smiling=yes) introduces changes to age, identity, and even adds glasses. These unintended shifts stem from global CFG treating all attributes equally. In contrast, DCFG applies decoupled guidance across attributes—assigning stronger weights to those affected by the intervention, allowing attributes that were not targeted by the intervention to remain unchanged. This results in counterfactuals that more faithfully reflect the intended change while preserving identity and consistency in non-intervened factors.
Figure 3: Reversibility analysis in CelebA-HQ $(64\times 64)$. Left: Quantitative evaluation of how well the original image is recovered after generating a counterfactual and mapping it back to the original condition under do(Smiling). Right: A qualitative example showing a counterfactual generated under do(Male) and its reconstruction after reversing the intervention with CFG and our DCFG.
Fig. 3 evaluates the reversibility of counterfactuals in CelebA-HQ. The left panel reports quantitative metrics (MAE and LPIPS) that assess how well the original image is recovered after applying an intervention (e.g., do(Smiling)) and then reversing it. With global CFG, reversibility consistently degrades as guidance strength increases, with higher error values across both metrics. In comparison, DCFG yields lower MAE and LPIPS for the same guidance levels on intervened variables (i.e., comparing $\omega$ with $\omega_{\text{aff}}$), demonstrating improved ability to recover the original image. The right panel presents a qualitative example where a counterfactual is generated under do(Male) and then reversed. Notably, the reversed image appears older than the original—an artifact not of the intervention itself, but of global CFG’s tendency to amplify all attributes in the counterfactual parent vector. While the model reverts the intervened attribute, non-intervened ones like Young = no remain amplified, leading to unintended and compounding changes. In contrast, the proposed DCFG distinguishes between intervened and invariant attributes, applying stronger guidance only to the target group. This mitigates attribute amplification in non-target attributes during both generation and reversal, enabling more faithful, disentangled, and reversible counterfactuals.
4.2 EMBED
We evaluate DCFG on the EMBED [18] breast mammography dataset. We define a binary circle attribute based on the presence of circular skin markers, and a binary breast density label, where categories A and B are grouped as low and categories C and D as high. See section E.1 for details.
Fig. 4 presents results for counterfactual generation on EMBED. The bar plot on the left reports $\Delta$ effectiveness metrics, measuring how classifier performance changes relative to the baseline. While global CFG improves effectiveness for the target attribute (circle), it also increases effectiveness on non-intervened attributes such as density, indicating unintended attribute amplfication. DCFG mitigates this by applying selective guidance, maintaining stable performance on non-target attributes. The figure on the right illustrates a key example: applying do(density) under global CFG unintentionally amplifies the presence of circular skin markers, as evidenced by the increased number of visible circles in both the counterfactual and reversed images. This is suppressed under DCFG, where circle features remain unchanged in both counterfactual and reversed images.
Figure 4: Evaluation of counterfactual generation on EMBED ($192\times 192$). Left: $\Delta$ metrics showing the effect of do(circle). DCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(density), the reversed image, and their difference maps (CF/Rev. - input). See section E.2 for full quantitative results and section E.3 for more visual results.
4.3 MIMIC
We evaluate our method on the MIMIC-CXR dataset [19]. We follow the dataset splits and filtering protocols from [9, 13], and focus on the binary disease label of pleural effusion. The underlying causal graph in De Sousa Ribeiro et al. [9] includes four attributes: race, sex, finding, and age. We adopt this setup, but since our goal is to study attribute amplification caused by CFG, we focus on sex, race, and finding, which we assume to be mutually independent for analysis purposes. See section F.1 for details.
Fig. 5 presents an evaluation of counterfactual generation in MIMIC-CXR, highlighting the advantages of our proposed DCFG. The bar plot on the left shows $\Delta$ metrics that quantify the change in effectiveness relative to the baseline $\omega=1.0$. While global CFG improves effectiveness for the intervened variable (finding), it also introduces substantial shifts in non-intervened attributes such as race and sex, revealing unwanted attribute amplfication. In contrast, DCFG achieves comparable or higher target effectiveness while suppressing spurious changes, demonstrating more precise and controlled generation. On the right, we show a qualitative example of a counterfactual generated under do(finding), its reversed reconstruction, and their corresponding difference maps. Compared to global CFG, our method yields localized, clinically meaningful changes in counterfactuals and better identity preservation in the reversed image—further supporting its robustness and causal faithfulness.
Figure 5: Evaluation of counterfactual generation on MIMIC ($192\times 192$). Left: $\Delta$ metrics showing the effect of do(finding). DCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(density), the reversed image, and their difference maps (CF/Rev. - input). See section F.2 for full quantitative results and section F.3 for more qualitative results.
5 Conclusion
We identified a limitation of standard CFG in counterfactual generation: it applies a uniform guidance weight to all conditioning attributes, leading to amplification of attributes that are meant to remain unchanged. To address this, we proposed Decoupled Classifier-Free Guidance (DCFG), a flexible framework that allows for arbitrary grouping of semantic attributes, with distinct guidance weights applied to each group during sampling, under the mild assumption that the groups are mutually independent. In this work, we instantiate DCFG for counterfactual image generation by separating attributes into two groups: intervened attributes, which include the intervened variables and their descendants in the causal graph, and invariant attributes, which are unaffected by the intervention and expected to remain unchanged. This structure enables DCFG to prevent spurious changes in invariant factors while preserving the intended effect of the intervention. While our focus is on counterfactuals, DCFG can be applied more broadly to conditional generation tasks that require group-wise control over conditioning signals. We evaluated DCFG on CelebA-HQ, EMBED, and MIMIC-CXR, covering natural and medical image domains. Our results show that DCFG significantly reduces attribute amplification while maintaining intervention effectiveness with improved identity preservation, particularly at higher guidance strengths.
Limitations and Future Work. While DCFG demonstrates promising results, the selection of $w_{\text{aff}}$ and $w_{\text{inv}}$ relies on empirical tuning. Future work could explore learned strategies to adaptively tune these weights based on the input condition or diffusion timestep. Additionally, DCFG’s modular nature opens up opportunities for broader integration. One direction is combining DCFG with latent diffusion models or diffusion autoencoders to enable high-resolution synthesis and stronger identity preservation. Another promising extension is applying DCFG selectively at specific diffusion timesteps. Prior work suggests that restricting guidance to mid-to-late timesteps can mitigate over-saturation and improve generation quality. Exploring such selective or dynamic scheduling strategies within the DCFG framework may further enhance counterfactual fidelity.
Acknowledgements
R.R. is supported by the Engineering and Physical Sciences Research Council (EPSRC) through a Doctoral Training Partnerships PhD Scholarship. A.K. is supported by UKRI (grant no. EP/S023356/1), as part of the UKRI Centre for Doctoral Training in Safe and Trusted AI. B.G. received support from the Royal Academy of Engineering as part of his Kheiron/RAEng Research Chair. B.G. and F.R. acknowledge the support of the UKRI AI programme, and the EPSRC, for CHAI - EPSRC Causality in Healthcare AI Hub (grant no. EP/Y028856/1).
Appendix A Background
A.1 Notation Summary
Symbol Description
$\mathbf{x}_{0}$ (also denoted as $\mathbf{x}$) Original observed image
$\mathbf{x}_{t}$ Noisy image at diffusion timestep $t$
$\mathbf{x}_{T}$ (i.e., $\mathbf{u}$) Latent code after DDIM forward process (abduction; exogenous noise)
$\tilde{\mathbf{x}}$ Generated counterfactual image
$\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\cdot)$ Denoiser prediction given condition input
$\mathbf{pa}$ Vector of semantic parent attributes (e.g., sex, age)
$\widetilde{\mathbf{pa}}$ Counterfactual parent attributes after intervention
$pa_{i}$ Raw value of the $i$-th semantic attribute
$\mathcal{E}_{i}$ Embedding MLP for $pa_{i}$: $\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d}$
$\mathbf{c}$ Full conditioning vector from $\mathbf{pa}$
$\tilde{\mathbf{c}}$ Conditioning vector from counterfactual attributes $\widetilde{\mathbf{pa}}$
$\varnothing$ Null token for classifier-free guidance (unconditional input)
$\omega_{m}$ CFG weight for attribute group $m$
$\mathbf{pa}^{(m)}$ Attributes in the $m$-th group
[…]
Table A.1: Notation used throughout the paper. Tilde ($\sim$) indicates counterfactual quantities.
A.2 Evaluating Counterfactuals
To evaluate the soundness of generated counterfactuals, we define a counterfactual image generation function $\mathcal{F}_{\theta}(\cdot)$, which produces counterfactuals according to
$\tilde{\mathbf{x}}:=\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}})=f_{\theta}(f_{\theta}^{-1}(\mathbf{x},\mathbf{pa}),\tilde{\mathbf{pa}}).$ (20)
We describe three key metrics used to assess counterfactual quality: composition, reversibility, and effectiveness [32, 30].
Composition evaluates how well the model reconstructs the original image under a null intervention, by computing a distance metric $d$ between the original image and its counterfactual:
$\text{Comp}(\mathbf{x},\mathbf{pa}):=d(\mathbf{x},\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\mathbf{pa})).$ (21)
Reversibility measures the consistency of the counterfactual transformation by applying the reverse intervention and comparing the result to the original image:
$\text{Rev}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}):=L_{1}(\mathbf{x},\mathcal{F}_{\theta}(\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}),\tilde{\mathbf{pa}},\mathbf{pa})).$ (22)
Effectiveness quantifies whether the intended intervention has the desired causal effect. It compares the intervened value $\tilde{\mathbf{pa}}_{k}$ with the prediction obtained by an anti-causal model applied to the counterfactual image:
$\text{Eff}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}):=L_{k}(\tilde{\mathbf{pa}}_{k},\mathbf{Pa}_{k}(\mathcal{F}_{\theta}(\mathbf{x},\mathbf{pa},\tilde{\mathbf{pa}}))).$ (23)
A note on Composition. We do not report the Composition metric in our evaluation, as it is ill-defined in the context of CFG and DCFG. Since both methods use the same trained diffusion model, applying a null intervention (i.e., $\tilde{\mathbf{pa}}=\mathbf{pa}$) does not meaningfully differentiate between them. If reconstruction is performed without guidance, CFG and DCFG are equivalent—reducing to standard decoding. If guidance is applied, it becomes unclear how to split attributes into invariant ($\mathbf{pa}_{\text{inv}}$) and intervened ($\mathbf{pa}_{\text{aff}}$) groups during null intervention. For instance, assigning all attributes to $\mathbf{pa}_{\text{inv}}$ with $\omega_{\text{inv}}=1$ would effectively disable guidance, making the comparison trivial and uninformative. For this reason, we focus on Effectiveness and Reversibility, which better capture the behavior of guided sampling under interventions.
Effectiveness Classifier. To evaluate effectiveness, we train a classifier with a ResNet-18 backbone for each dataset, using the split as in table A.2. The classifier predicts the intervened attribute from generated counterfactuals, and AUROC is used to quantify intervention success. On CelebA, the classifier achieves AUROC scores of 0.974 (Smiling), 0.992 (Male), and 0.828 (Young). On EMBED, the AUROC is 0.935 for density and 0.908 for circle. On MIMIC-CXR, AUROC scores are 0.864 for race, 0.991 for sex, and 0.938 for finding.
Reversibility Metrics. We use Mean Absolute Error (MAE) and LPIPS [69] to evaluate reversibility. These metrics quantify the pixel-level and perceptual similarity, respectively, between the original image and its reversed counterfactual.
Appendix B Decoupled Classifier-Free Guidance
We provide a theoretical justification for the Decoupled Classifier-Free Guidance formulation presented in Proposition 1, interpreting it as gradient ascent on a sharpened proxy posterior under a group-level conditional independence assumption.
Proof of Proposition 1.
We begin by assuming that the semantic attributes are partitioned into $M$ disjoint groups:
$\mathbf{pa}=(\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}).$ (24)
Under the assumption that these groups are conditionally independent given $\mathbf{x}_{t}$, we have:
$p(\mathbf{pa}\mid\mathbf{x}_{t})=\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t}).$ (25)
Applying Bayes’ rule:
$p(\mathbf{x}_{t}\mid\mathbf{pa})=\frac{p(\mathbf{pa}\mid\mathbf{x}_{t})\cdot p(\mathbf{x}_{t})}{p(\mathbf{pa})}\propto p(\mathbf{pa}\mid\mathbf{x}_{t})\cdot p(\mathbf{x}_{t}).$ (26)
so the posterior can be factorized as:
$p(\mathbf{x}_{t}\mid\mathbf{pa})\propto p(\mathbf{x}_{t})\cdot\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t}).$ (27)
Applying group-level guidance weights $\omega_{m}$ yields the sharpened proxy posterior:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})\propto p(\mathbf{x}_{t})\cdot\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})^{\omega_{m}}.$ (28)
Gradient of the Log Proxy Posterior.
For DCFG, the gradient becomes:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})$ $=\nabla\log p(\mathbf{x}_{t})+\sum_{m=1}^{M}\omega_{m}\cdot\left(\nabla\log p(\mathbf{x}_{t}\mid\mathbf{pa}^{(m)})-\nabla\log p(\mathbf{x}_{t})\right),$ (29)
where $\mathbf{pa}^{(m)}$ is the $m$-th group of attributes, and $\omega_{m}$ is the guidance weight for that group.
The corresponding implementation interpolates denoising scores per group:
[…]
Appendix C Implementation Details
Architecture
We adopt the commonly used U-net backbone [10] for all diffusion models in this work. We modify it to support CFG and DCFG. Each conditioning attribute is projected via a dedicated MLP encoder (section 3.1), and the resulting embeddings are concatenated with the timestep embedding. During training, we apply exponential moving average (EMA) to model weights for improved stability. All images are normalized to the range [-1,1]. The complete architecture and training configurations for each dataset are summarized in table A.2.
Table A.2: Training and architecture of diffusion U-Net configurations used in our experiments. *We did not evaluate on the whole test set due to the high computational cost of diffusion sampling.
Parameter CelebA-HQ EMBED MIMIC-CXR
Train Set Size 24,000 151,948 62,336
Validation Set Size 3,000 7,156 9,968
Test Set Size 3,000 43,669 30,535
Resolution $64\times 64\times 3$ $192\times 192\times 1$ $192\times 192\times 1$
Batch Size 128 48 48
Training Epochs 6000 5000 5000
Base Channels (U-Net) 64 64 32
Channel Multipliers [1,2,4,8] [1,1,2,2,4,4] [1,2,3,4,5,6]
Attention Resolutions [16] - -
ResNet Blocks 2 2 2
Dropout Rate 0.1 0.0 0.0
Num. Conditioning Attrs 3 2 4
Cond. Embedding Dim $3\times 96$ $2\times 64$ $4\times 64$
Noise Schedule Linear Cosine Linear
Learning Rate 1e-4
Optimiser Adam (wd 1e-4)
EMA Decay 0.9999
Training Steps $T$ 1000
Loss MSE (noise prediction)
Training Procedure for DCFG
The training of our proposed DCFG follows the same setup as standard Classifier-Free Guidance (CFG). Specifically, we apply classifier-free dropout [16] by replacing the entire conditioning vector with a null token (i.e., zero vector) with probability $p_{\varnothing}=0.5$. Unlike the typical choice of $p_{\varnothing}=0.2$, we found that using $p_{\varnothing}=0.5$ better preserves identity, which is particularly important for counterfactual generation tasks. Note that we apply dropout to all attributes jointly, rather than selectively masking subsets. One could alternatively consider group-wise dropout—nullifying only a random subset of attribute groups—but such partial masking may encourage the model to over-rely on the remaining visible attributes, making the resulting guidance less disentangled and less robust. We leave this as an interesting direction for future exploration.
Computation Resources
All experiments were conducted on servers equipped with multiple NVIDIA GPUs, including L40S and similar models, each with approximately 48GB of memory. Training each model typically takes around one week on one GPU. Due to the high computational cost of diffusion-based sampling, generating counterfactuals for each intervention (do(key)) and each guidance configuration takes approximately one day for the MIMIC-CXR and EMBED datasets, and around 7 hours for CelebA-HQ.
Evaluation
Due to the computational cost of diffusion sampling, we evaluate on fixed, balanced subsets rather than the full test sets. For CelebA-HQ and EMBED, we use 1,000 samples each, selected to ensure an even distribution across the conditioning attributes. For MIMIC-CXR, we evaluate on 1,500 samples, stratified to balance race groups. These fixed subsets are reused across all experiments to enable fair and consistent comparisons. To generate counterfactuals, we use DDIM sampling with 1,000 time steps, as we find this setting achieves stronger identity preservation compared to shorter schedules—an essential property for counterfactual evaluation. This setup allows us to assess attribute-specific phenomena such as amplification and reversibility while keeping the sampling cost manageable.
Appendix D CelebA-HQ
D.1 Dataset details
For the CelebA-HQ dataset [20], we select Smiling, Male, and Young as three independent binary variables. These are among the most reliably annotated attributes in CelebA, each achieving over 95% consistency in manual labeling [60]. Moreover, they exhibit low inconsistency across duplicate face pairs (e.g., Male: 0.005; Smiling: 0.077), suggesting minimal label noise. We assume these variables to be independent, as our goal is to isolate and analyze attribute amplification under global classifier-free guidance (CFG), which is more interpretable with uncorrelated factors. As shown in Fig. A.1, the Pearson correlation matrix confirms weak pairwise correlations among these attributes. Although a moderate negative correlation is observed between Male and Young ($\rho=-0.33$), we attribute this to dataset bias rather than a true causal dependency, and proceed by modeling them as independent.
Figure A.1: Pearson correlation matrix of CelebA-HQ attributes: Smiling, Male, and Young. While a moderate negative correlation is observed between Male and Young ($\rho=-0.33$), we regard this as a spurious correlation likely stemming from dataset bias rather than a meaningful causal relationship. Therefore, for the purposes of our analysis, we assume these attributes to be independent.
D.2 Extra quantitative result for Celeba-HQ
Table A.3: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility deteriorates with increasing $\omega_{\mathrm{aff}}$, DCFG consistently maintains better reversibility than global CFG with $\omega=\omega_{\mathrm{aff}}$.
do(key) CFG Smiling AUC/$\Delta$ Male AUC/$\Delta$ Young AUC/$\Delta$ MAE LPIPS
do(Smiling) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
$\omega=1.2$ 91.7 / +5.2 98.5 / +1.6 80.9 / +2.3 0.133 0.091
$\omega=1.5$ 95.5 / +9.0 99.2 / +2.3 84.3 / +5.7 0.163 0.111
$\omega=1.7$ 96.7 / +10.2 99.4 / +2.5 85.7 / +7.1 0.179 0.119
$\omega=2.0$ 97.7 / +11.2 99.6 / +2.7 88.3 / +9.7 0.203 0.127
$\omega=2.5$ 98.6 / +12.1 99.7 / +2.8 89.3 / +10.7 0.234 0.142
$\omega=3.0$ 98.8 / +12.3 99.9 / +3.0 90.3 / +11.7 0.263 0.155
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 89.9 / +3.4 95.8 / -1.1 77.8 / -0.8 0.128 0.093
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.2$ 93.9 / +7.4 97.5 / +0.6 79.7 / +1.1 0.136 0.092
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 95.4 / +8.9 97.5 / +0.6 79.9 / +1.3 0.141 0.095
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 97.0 / +10.5 97.3 / +0.4 79.7 / +1.1 0.146 0.098
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 98.9 / +12.4 96.1 / -0.8 77.8 / -0.8 0.164 0.112
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 99.6 / +13.1 95.4 / -1.5 77.6 / -1.0 0.177 0.122
do(Male) $\omega=1.0$ 86.6 / +0.0 91.8 / +0.0 79.8 / +0.0 0.115 0.079
$\omega=1.2$ 90.1 / +3.5 95.1 / +3.3 80.8 / +1.0 0.127 0.088
$\omega=1.5$ 93.3 / +6.7 97.2 / +5.4 82.0 / +2.2 0.158 0.111
$\omega=1.7$ 94.7 / +8.1 97.5 / +5.7 83.7 / +3.9 0.175 0.123
$\omega=2.0$ 96.0 / +9.4 97.9 / +6.1 85.0 / +5.2 0.202 0.139
$\omega=2.5$ 97.6 / +11.0 98.4 / +6.6 87.5 / +7.7 0.238 0.156
$\omega=3.0$ 98.2 / +11.6 99.2 / +7.4 90.2 / +10.4 0.267 0.171
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 85.1 / -1.5 91.3 / -0.5 79.0 / -0.8 0.137 0.097
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.2$ 88.3 / +1.7 93.8 / +2.0 78.9 / -0.9 0.149 0.101
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 88.1 / +1.5 95.8 / +4.0 77.5 / -2.3 0.151 0.103
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 88.0 / +1.4 97.8 / +6.0 77.0 / -2.8 0.158 0.109
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 87.4 / +0.8 99.4 / +7.6 76.2 / -3.6 0.171 0.118
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 87.4 / +0.8 99.7 / +7.9 75.9 / -3.9 0.188 0.130
do(Young) $\omega=1.0$ 87.5 / +0.0 95.7 / +0.0 62.3 / +0.0 0.115 0.085
$\omega=1.2$ 90.8 / +3.3 97.4 / +1.7 64.5 / +2.2 0.130 0.088
$\omega=1.5$ 95.6 / +8.1 99.3 / +3.6 66.3 / +4.0 0.166 0.110
$\omega=1.7$ 96.7 / +9.2 99.4 / +3.7 67.8 / +5.5 0.183 0.119
$\omega=2.0$ 97.7 / +10.2 99.6 / +3.9 69.5 / +7.2 0.204 0.130
$\omega=2.5$ 98.3 / +10.8 99.8 / +4.1 73.5 / +11.2 0.234 0.146
$\omega=3.0$ 98.5 / +11.0 99.9 / +4.2 77.7 / +15.4 0.261 0.160
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 87.4 / -0.1 95.0 / -0.7 63.2 / +0.9 0.129 0.095
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.2$ 90.0 / +2.5 96.7 / +1.0 67.4 / +5.1 0.147 0.100
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 89.2 / +1.7 96.1 / +0.4 71.3 / +9.0 0.150 0.103
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 87.9 / +0.4 94.5 / -1.2 75.6 / +13.3 0.157 0.110
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 88.5 / +1.0 91.9 / -3.8 81.8 / +19.5 0.172 0.125
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 86.7 / -0.8 90.0 / -5.7 87.6 / +25.3 0.188 0.136
Table A.4: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Smiling AUC/$\Delta$ Male AUC/$\Delta$ Young AUC/$\Delta$ MAE LPIPS
do(Smiling) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
$\omega=2.5$ 98.6 / +12.1 99.7 / +2.8 89.3 / +10.7 0.234 0.142
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 99.1 / +12.6 92.6 / -4.3 75.2 / -3.4 0.165 0.118
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 98.9 / +12.4 96.1 / -0.8 77.8 / -0.8 0.164 0.112
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 98.3 / +11.8 98.2 / +1.3 82.6 / +4.0 0.177 0.118
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 98.1 / +11.6 98.8 / +1.9 84.0 / +5.4 0.189 0.123
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 97.5 / +11.0 99.3 / +2.4 87.0 / +8.4 0.209 0.131
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 96.3 / +9.8 99.5 / +2.6 88.7 / +10.1 0.236 0.143
do(Male) $\omega=1.0$ 86.6 / +0.0 91.8 / +0.0 79.8 / +0.0 0.115 0.079
$\omega=2.5$ 97.6 / +11.0 98.4 / +6.6 87.5 / +7.7 0.238 0.156
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 83.4 / -3.2 99.4 / +7.6 68.9 / -10.9 0.173 0.122
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 87.4 / +0.8 99.4 / +7.6 71.1 / -8.7 0.171 0.118
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 92.0 / +5.4 99.3 / +7.5 74.0 / -5.8 0.182 0.119
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 93.5 / +6.9 98.9 / +7.1 76.8 / -3.0 0.189 0.123
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 95.3 / +8.7 98.7 / +6.9 80.2 / +0.4 0.207 0.135
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 97.2 / +10.6 97.7 / +5.9 87.5 / +7.7 0.242 0.158
do(Young) $\omega=1.0$ 87.5 / +0.0 95.7 / +0.0 62.3 / +0.0 0.115 0.085
$\omega=2.5$ 98.3 / +10.8 99.8 / +4.1 73.5 / +11.2 0.234 0.146
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 83.4 / -4.1 85.9 / -9.8 85.1 / +22.8 0.169 0.127
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 88.5 / +1.0 91.9 / -3.8 81.8 / +19.5 0.172 0.125
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 92.4 / +4.9 96.4 / +0.7 78.5 / +16.2 0.187 0.127
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 94.3 / +6.8 97.8 / +2.1 75.3 / +13.0 0.199 0.133
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 97.1 / +9.6 99.0 / +3.3 73.3 / +11.0 0.215 0.139
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 99.3 / +11.8 99.7 / +4.0 68.5 / +6.2 0.238 0.147
D.3 Extra qualitative results for CelebA-HQ
Figure A.2: Additional qualitative results for do(Male) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves non-intervened attributes and identity while effectively reflecting the intervention.
Figure A.3: Additional qualitative results for do(Smiling) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
Figure A.4: Additional qualitative results for do(Young) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
Figure A.5: Reversibility analysis for do(Similing) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
Figure A.6: Reversibility analysis for do(Male) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
Figure A.7: Reversibility analysis for do(Young) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
Figure A.8: Effect of $\omega_{\text{inv}}$ on CelebA-HQ counterfactuals. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega=2.5$) and our proposed DCFG with fixed intervention guidance ($\omega_{\text{aff}}=2.5$) and varying invariant guidance $\omega_{\text{inv}}\in\{1.0,1.5,2.0,2.5\}$. As $\omega_{\text{inv}}$ increases, amplification of invariant attributes becomes more pronounced, and at $\omega_{\text{inv}}=2.5$, DCFG effectively reproduces the same over-editing behavior as global CFG. This shows that $\omega_{\text{inv}}$ modulates the degree of guidance applied to invariant attributes and should be carefully calibrated to maintain identity and disentanglement.
Appendix E EMEBD
E.1 Dataset details
We use the EMory BrEast imaging Dataset (EMBED) [18] for our experiments. Schueppert et al. [48] manually labeled 22,012 images with circular markers and trained a classifier on this subset, which was then applied to the full dataset to infer circle annotations. We adopt this preprocessing pipeline and extract the circle attribute from their predictions. To define the density label, we binarize the original four-category breast density annotations by grouping categories A and B as low density, and categories C and D as high density. While the full dataset comprises 151,948 training, 7,156 validation, and 43,669 test samples, we use only 1,000 test samples in this work due to the high computational cost of diffusion models. As shown in Fig. A.9, the Pearson correlation matrix reveals that density and circle are nearly uncorrelated, supporting our assumption of their independence.
Figure A.9: Pearson correlation matrix of EMBED attributes: density and circle. The correlation between these two variables is negligible ($\rho=-0.04$), suggesting that they can be reasonably treated as independent for the purposes of our analysis.
E.2 Extra quantitative results for EMBED
Table A.5: EMBED: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility deteriorates with increasing $\omega_{\mathrm{aff}}$, DCFG consistently maintains better reversibility than global CFG with $\omega=\omega_{\mathrm{aff}}$.
do(key) CFG Density AUC/$\Delta$ Circle AUC/$\Delta$ MAE LPIPS
do(density) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
$\omega=1.2$ 70.5 / +7.1 94.5 / +1.6 0.033 0.038
$\omega=1.5$ 79.0 / +15.6 95.9 / +3.0 0.035 0.047
$\omega=1.7$ 84.3 / +20.9 96.7 / +3.8 0.032 0.055
$\omega=2.0$ 89.6 / +26.2 97.5 / +4.6 0.034 0.064
$\omega=2.5$ 95.2 / +31.8 97.7 / +4.8 0.042 0.076
$\omega=3.0$ 97.8 / +34.4 98.2 / +5.3 0.045 0.086
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 73.1 / +9.7 92.8 / -0.1 0.028 0.038
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.0$ 81.6 / +18.2 92.2 / -0.7 0.029 0.043
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.0$ 86.2 / +22.8 91.6 / -1.3 0.031 0.048
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.0$ 91.6 / +28.2 90.7 / -2.2 0.032 0.053
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 96.6 / +33.2 92.2 / -0.7 0.036 0.064
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 98.6 / +35.2 91.6 / -1.3 0.038 0.071
do(circle) $\omega=1.0$ 92.6 / +0.0 90.6 / +0.0 0.023 0.026
$\omega=1.2$ 94.7 / +2.1 92.1 / +1.5 0.029 0.024
$\omega=1.5$ 96.8 / +4.2 93.9 / +3.3 0.030 0.027
$\omega=1.7$ 97.9 / +5.3 95.2 / +4.6 0.027 0.035
$\omega=2.0$ 98.8 / +6.2 96.4 / +5.8 0.030 0.040
$\omega=2.5$ 99.7 / +7.1 97.8 / +7.2 0.038 0.043
$\omega=3.0$ 99.9 / +7.3 98.4 / +7.8 0.042 0.051
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 93.3 / +0.7 92.6 / +2.0 0.024 0.028
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.0$ 93.2 / +0.6 94.4 / +3.8 0.025 0.030
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.0$ 93.2 / +0.6 95.7 / +5.1 0.025 0.032
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.0$ 92.9 / +0.3 97.2 / +6.6 0.026 0.034
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 94.5 / +1.9 98.5 / +7.9 0.027 0.038
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 94.0 / +1.4 98.9 / +8.3 0.029 0.042
Table A.6: EMBED: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Density AUC/$\Delta$ Circle AUC/$\Delta$ MAE LPIPS
do(density) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
$\omega=2.5$ 95.2 / +31.8 97.7 / +4.8 0.042 0.076
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 96.7 / +33.3 89.5 / -3.4 0.035 0.063
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 96.6 / +33.2 92.2 / -0.7 0.036 0.064
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 96.6 / +33.2 94.6 / +1.7 0.036 0.067
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 96.6 / +33.2 95.7 / +2.8 0.037 0.070
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 96.5 / +33.1 96.6 / +3.7 0.038 0.073
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 96.4 / +33.0 97.6 / +4.7 0.039 0.080
do(circle) $\omega=1.0$ 92.6 / +0.0 90.6 / +0.0 0.023 0.026
$\omega=2.5$ 99.7 / +7.1 97.8 / +7.2 0.038 0.043
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 92.2 / -0.4 98.5 / +7.9 0.028 0.039
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 94.5 / +1.9 98.5 / +7.9 0.027 0.038
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 97.2 / +4.6 98.3 / +7.7 0.028 0.039
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 98.1 / +5.5 98.2 / +7.6 0.029 0.040
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 99.0 / +6.4 98.2 / +7.6 0.031 0.043
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 99.8 / +7.2 98.0 / +7.4 0.035 0.050
E.3 Extra qualitative results for EMBED
Figure A.10: Additional qualitative results for do(density) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in circle count, whereas DCFG mitigates such amplification.
Figure A.11: Additional qualitative results for do(circle) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$) and the difference map (CF-input). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in density, whereas DCFG mitigates such amplification.
Figure A.12: Reversibility analysis for do(density) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). DCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Figure A.13: Reversibility analysis for do(circle) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). DCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Appendix F MIMIC
F.1 Dataset details
We use the MIMIC-CXR dataset [19] in our experiments. Following the dataset splits and filtering protocols of [9, 13], we focus on the binary disease label for pleural effusion. We adopt the same causal graph (DAG) as proposed in De Sousa Ribeiro et al. [9], in which age is modeled as a parent of finding. While we include age as part of the conditioning variables, we do not intervene on it. Instead, our primary goal is to study amplification of unintervened variables caused by CFG. For this purpose, we focus on race, sex, and finding, which we assume to be mutually independent. Fig. A.14 shows the Pearson correlation matrix of these three attributes, where all pairwise correlations are small (e.g., $\rho=0.12$ between race and sex, and $\rho=-0.15$ between race and finding), supporting the validity of the independence assumption in our counterfactual modeling.
Figure A.14: Pearson correlation matrix of MIMIC attributes: race, sex, and finding. All pairwise correlations are low (e.g., $\rho=0.12$ between race and sex, and $\rho=-0.15$ between race and finding), suggesting that these variables can be reasonably treated as independent for the purposes of our counterfactual analysis.
F.2 Extra quantitative results for MIMIC-CXR
Table A.7: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility tends to degrade as $\omega_{\mathrm{aff}}$ increases, DCFG maintains better reversibility than global CFG at higher guidance strengths.
do(key) CFG Sex AUC/$\Delta$ Race AUC/$\Delta$ Finding AUC/$\Delta$ MAE LPIPS
do(sex) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
$\omega=1.2$ 95.2 / +2.8 79.3 / +3.7 92.8 / +4.0 0.151 0.206
$\omega=1.5$ 97.7 / +5.3 82.4 / +6.8 95.7 / +6.9 0.171 0.226
$\omega=1.7$ 98.5 / +6.1 84.7 / +9.1 97.0 / +8.2 0.186 0.239
$\omega=2.0$ 99.3 / +6.9 87.4 / +11.8 98.0 / +9.2 0.207 0.258
$\omega=2.5$ 99.8 / +7.4 90.5 / +14.9 99.0 / +10.2 0.239 0.284
$\omega=3.0$ 99.9 / +7.5 92.9 / +17.3 99.5 / +10.7 0.266 0.305
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 96.4 / +4.0 74.6 / -1.0 89.1 / +0.3 0.158 0.217
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.0$ 98.4 / +6.0 74.1 / -1.5 88.5 / -0.3 0.167 0.227
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 99.2 / +6.8 76.9 / +1.3 91.5 / +2.7 0.174 0.233
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 99.5 / +7.1 75.8 / +0.2 90.9 / +2.1 0.183 0.243
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 99.9 / +7.5 74.9 / -0.7 90.1 / +1.3 0.199 0.260
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 100.0 / +7.6 74.5 / -1.1 89.4 / +0.6 0.216 0.276
do(race) $\omega=1.0$ 95.1 / +0.0 65.4 / +0.0 90.4 / +0.0 0.135 0.191
$\omega=1.2$ 97.6 / +2.5 69.9 / +4.5 93.3 / +2.9 0.135 0.190
$\omega=1.5$ 98.9 / +3.8 73.9 / +8.5 96.2 / +5.8 0.155 0.209
$\omega=1.7$ 99.3 / +4.2 76.3 / +10.9 97.5 / +7.1 0.171 0.223
$\omega=2.0$ 99.6 / +4.5 80.5 / +15.1 98.2 / +7.8 0.193 0.242
$\omega=2.5$ 99.7 / +4.6 86.1 / +20.7 99.2 / +8.8 0.229 0.271
$\omega=3.0$ 99.8 / +4.7 90.1 / +24.7 99.4 / +9.0 0.256 0.292
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 95.4 / +0.3 69.6 / +4.2 90.3 / -0.1 0.141 0.198
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.0$ 94.8 / -0.3 75.5 / +10.1 90.0 / -0.4 0.147 0.203
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 97.2 / +2.1 78.3 / +12.9 92.4 / +2.0 0.153 0.208
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 96.4 / +1.3 82.8 / +17.4 92.0 / +1.6 0.160 0.215
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 95.6 / +0.5 89.0 / +23.6 91.7 / +1.3 0.178 0.231
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 94.0 / -1.1 92.7 / +27.3 91.7 / +1.3 0.197 0.249
do(finding) $\omega=1.0$ 94.6 / +0.0 78.3 / +0.0 80.8 / +0.0 0.134 0.193
$\omega=1.2$ 97.2 / +2.6 81.2 / +2.9 85.7 / +4.9 0.136 0.194
$\omega=1.5$ 98.9 / +4.3 83.8 / +5.5 90.6 / +9.8 0.153 0.210
$\omega=1.7$ 99.5 / +4.9 85.7 / +7.4 92.9 / +12.1 0.165 0.222
$\omega=2.0$ 99.7 / +5.1 88.5 / +10.2 95.0 / +14.2 0.184 0.239
$\omega=2.5$ 99.8 / +5.2 91.9 / +13.6 97.7 / +16.9 0.215 0.264
$\omega=3.0$ 99.9 / +5.3 93.8 / +15.5 98.6 / +17.8 0.244 0.287
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 95.1 / +0.5 78.1 / -0.2 85.6 / +4.8 0.142 0.202
$\omega_{\mathrm{aff}}=1.5$, $\omega_{\mathrm{inv}}=1.0$ 94.9 / +0.3 77.8 / -0.5 92.0 / +11.2 0.141 0.201
$\omega_{\mathrm{aff}}=1.7$, $\omega_{\mathrm{inv}}=1.2$ 97.1 / +2.5 80.6 / +2.3 93.5 / +12.7 0.150 0.209
$\omega_{\mathrm{aff}}=2.0$, $\omega_{\mathrm{inv}}=1.2$ 96.9 / +2.3 80.2 / +1.9 96.6 / +15.8 0.149 0.209
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 96.4 / +1.8 80.1 / +1.8 98.8 / +18.0 0.151 0.212
$\omega_{\mathrm{aff}}=3.0$, $\omega_{\mathrm{inv}}=1.2$ 95.2 / +0.6 79.3 / +1.0 99.6 / +18.8 0.154 0.216
Table A.8: MIMIC: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Sex AUC/$\Delta$ Race AUC/$\Delta$ Finding AUC/$\Delta$ MAE LPIPS
do(sex) $\omega=1.0$ 92.4 / +0.0 75.6 / +0.0 88.8 / +0.0 0.146 0.202
$\omega=2.5$ 99.8 / +7.4 90.5 / +14.9 99.0 / +10.2 0.239 0.284
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 99.9 / +7.5 71.3 / -4.3 86.2 / -2.6 0.200 0.261
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 99.9 / +7.5 74.9 / -0.7 90.1 / +1.3 0.199 0.260
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 99.8 / +7.4 80.1 / +4.5 94.2 / +5.4 0.207 0.264
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 99.7 / +7.3 83.2 / +7.6 95.9 / +7.1 0.214 0.269
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 99.7 / +7.3 86.7 / +11.1 97.5 / +8.7 0.227 0.278
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 99.6 / +7.2 90.3 / +14.7 98.9 / +10.1 0.249 0.293
do(race) $\omega=1.0$ 95.1 / +0.0 65.4 / +0.0 90.4 / +0.0 0.135 0.191
$\omega=2.5$ 99.7 / +4.6 86.1 / +20.7 99.2 / +8.8 0.229 0.271
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 91.3 / -3.8 89.5 / +24.1 88.4 / -2.0 0.181 0.237
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 95.6 / +0.5 89.0 / +23.6 91.7 / +1.3 0.178 0.231
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 98.5 / +3.4 87.9 / +22.5 95.2 / +4.8 0.185 0.236
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 99.2 / +4.1 87.4 / +22.0 96.8 / +6.4 0.191 0.242
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 99.6 / +4.5 86.3 / +20.9 98.0 / +7.6 0.205 0.253
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 99.8 / +4.7 85.6 / +20.2 99.1 / +8.7 0.231 0.274
do(finding) $\omega=1.0$ 94.6 / +0.0 78.3 / +0.0 80.8 / +0.0 0.134 0.193
$\omega=2.5$ 99.8 / +5.2 91.9 / +13.6 97.7 / +16.9 0.215 0.264
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 93.0 / -1.6 77.0 / -1.3 99.0 / +18.2 0.143 0.206
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.2$ 96.4 / +1.8 80.1 / +1.8 98.8 / +18.0 0.151 0.212
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.5$ 98.5 / +3.9 84.1 / +5.8 98.3 / +17.5 0.166 0.224
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.7$ 99.2 / +4.6 86.2 / +7.9 97.8 / +17.0 0.180 0.236
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.0$ 99.6 / +5.0 89.4 / +11.1 97.1 / +16.3 0.202 0.254
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=2.5$ 99.9 / +5.3 92.6 / +14.3 95.8 / +15.0 0.239 0.282
F.3 Extra visual results for MIMIC
Figure A.15: Additional qualitative results for do(finding) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while accurately applying the intended intervention. Compared to standard CFG, DCFG produces counterfactuals with more localized changes and stronger identity preservation.
Figure A.16: Additional qualitative results for do(sex) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while accurately applying the intended intervention on sex. Compared to standard CFG, which tends to amplify unrelated features such as disease (i.e. finding), DCFG produces counterfactuals with more localized, semantically aligned changes and stronger identity preservation.
Figure A.17: Additional qualitative results for do(race) on MIMIC. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). While race interventions correspond to relatively subtle visual changes, standard CFG often amplifies unrelated features such as disease appearance (e.g., finding). In contrast, DCFG better preserves invariant attributes and identity, producing counterfactuals that are more localized, semantically aligned, and faithful to the intended intervention.
Figure A.18: Reversibility analysis on MIMIC. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). DCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Appendix G Broader Impacts
This work introduces a method for more faithful and disentangled counterfactual generation using diffusion models, which can benefit applications in healthcare, fairness auditing, and scientific analysis by enabling clearer causal reasoning and more interpretable image edits. However, as with other generative models, it also poses risks of misuse, such as enabling more precise deepfakes or reinforcing dataset biases. We recommend applying appropriate safeguards and fairness evaluations when deploying such methods in real-world settings.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the main claims of the paper, including the introduction of DCFG for targeted image intervention, its advantages over global CFG, and empirical findings regarding effectiveness and amplification. The motivation is clearly presented, and the claims made are consistent with the theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a "Limitations and Future Work" paragraph inside section 5 that explicitly discusses the reliance on empirically tuned hyperparameters ($\omega_{\text{aff}}$, $\omega_{\text{inv}}$), and suggests potential improvements through adaptive tuning and selective timestep scheduling. It also reflects on the scope of the current experiments and outlines directions for extending DCFG to high-resolution models and broader use cases, thereby transparently acknowledging the method’s assumptions and areas for improvement.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See proposition 1, and full proof for proposition 1 is in appendix B
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe the method clearly in section 3, and we provide implementation details in appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Data are publicly available. Code is not publicly available at the moment, but will be released later.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Our effectiveness results are reported in terms of AUROC. However, due to the high computational cost associated with running multiple training seeds or data splits, we did not report standard deviation or confidence intervals.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See appendix C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We read the Code of Ethics and make sure we obey it.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See appendix G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: N/A. This paper is not releasing any new dataset.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Data are publicly available, and original papers introducing these data are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper is not releasing any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The focus of this paper is not relevant to LLM.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.